OpenStudy (anonymous):

WILL FAN AND MEDAL!!~~ the square root of 169^5

OpenStudy (anonymous):

\[\sqrt[5]{169}\]

OpenStudy (anonymous):

is the equation like this?

OpenStudy (anonymous):

371293

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\[\sqrt{169^{5}}\]

OpenStudy (anonymous):

371293

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

\[\Large \sqrt{169^5} = 169^{\frac{5}{2}} = 371293\]

OpenStudy (anonymous):

371293 for the second 1

OpenStudy (anonymous):

was i right

OpenStudy (anonymous):

Oh I messed up the square root of 169x^5 @johnweldon1993

OpenStudy (johnweldon1993):

That's what I figured it was lol \[\large \sqrt{169x^5}\]

OpenStudy (johnweldon1993):

Hint* what is the square root of 169?

OpenStudy (anonymous):

gee thanks haha so how do i do it?

OpenStudy (johnweldon1993):

We know the square root of 169 = 13....so we have \[\large 13\sqrt{x^5}\] Now...what else can we do?

OpenStudy (johnweldon1993):

*hint* x^5 can be written as x^4 * x

OpenStudy (anonymous):

is the final answer 609.3?

OpenStudy (anonymous):

Your message got me so lost. @johnweldon1993